# Minquan Cheng, Huimei Wei, Kai Wan, and Giuseppe Caire

### v2

@@ Abstract
Coded caching is a promising technique to effectively reduce peak traffic by using local caches and the multicast gains generated by these local caches. We prefer to design a coded caching scheme with the subpacketization $F$ and transmission load $R$ as small as possible since these are the key metrics for evaluating the implementation complexity and transmission efficiency of the scheme, respectively. However, most of the existing coded caching schemes have large subpacketizations which grow exponentially with the number of users $K$, and there are a few schemes with linear subpacketizations which have large transmission loads. In this paper, we focus on studying the linear subpacketization, i.e., $K=F$, coded caching scheme with low transmission load. Specifically, we first introduce a new combinatorial structure called non-half-sum disjoint packing (NHSDP) which can be used to generate a coded caching scheme with $K=F$. Then a class of new schemes is obtained by constructing NHSDP. Theoretical and numerical comparisons show that (i) compared to the existing schemes with linear subpacketization (to the number of users), the proposed scheme achieves a lower load; (ii) compared to some existing schemes with polynomial subpacketization, the proposed scheme can also achieve a lower load in some cases; (iii) compared to some existing schemes with exponential subpacketization, the proposed scheme has loads close to those of these schemes in some cases. Moreover, the new concept of NHSDP is closely related to the classical combinatorial structures such as cyclic difference packing (CDP), non-three-term arithmetic progressions (NTAP), and perfect hash family (PHF). These connections indicate that NHSDP is an important combinatorial structure in the field of combinatorial design.
@@
Coded caching scheme, placement delivery array, linear subpacketization, transmission load, non-half-sum disjoint packing
@@ I INTRODUCTION
The rise of wireless devices in recent years has significantly increased network traffic, fueled by activities such as multimedia streaming, web browsing, and social networking. In addition, the high temporal variability of this traffic leads to congestion during peak periods and inefficient use of network resources during off-peak times. Caching is a promising technique to reduce peak traffic by taking advantage of memories distributed across the network to duplicate content during off-peak times. Conventional uncoded caching techniques focus on predicting the user demands for making an appropriate prefetching strategy, thus realizing a “local caching gain”, which scales with the amount of local memory [1]. In their seminal paper [2], Maddah-Ali and Niesen (MN) demonstrated that, in addition to a local caching gain, coded caching can also attain a “global caching gain”. This global caching gain scales with the global amount of memory in the network, since each transmission of the MN scheme can serve multiple users simultaneously.
@@
In a $(K,M,N)$ coded caching system, there is a central server having $N$ equal-sized files and $K$ users each of whom can cache up to $M$ files. The server connects to the users over an error-free shared link. An $F$-division $(K,M,N)$ coded caching scheme consists of two phases: the placement phase during the off-peak hours and the delivery phase during the peak hours. In the placement phase, the server divides each file into $F$ equal packets and places some of these packets into each user’s cache without knowledge of the users’ future demands. If packets are directly placed in each user’s cache, it is referred to as uncoded placement. Otherwise it is called coded placement. The parameter $F$ is called the subpacketization. In the delivery phase, each user requests one file randomly. According to the users’ demands and the caching packets, the server broadcasts some coded packets such that each user can recover its desired file. The normalized amount of transmission for the worst-case over all possible demands is called the transmission load (or load) $R$.
@@
The first well-known coded caching scheme was proposed by Maddah-Ali and Niesen in [2], which utilized combinatorial design in the placement phase and linear coding in the delivery phase. This scheme is referred to as the MN scheme. The authors in [3, 4, 5, 6] separately proved the minimality of the load of the MN scheme under the uncoded placement and for the case $K\leq N$. Using the placement strategy of the MN scheme, the authors in [5] proposed an exact optimal caching scheme, which strictly improves upon the current state of the art by exploiting commonalities among user demands under uncoded placement.
@@
However, the subpacketization $F={K\choose KM/N}$ of the MN scheme increases exponentially with the number of users $K$. This would become infeasible for practical implementation when $K$ is large. So it is meaningful to design a coded caching scheme with low subpacketization. The grouping method, which is widely regarded as the most effective in reducing subpacketization, was proposed in [7, 8]. Nevertheless, the load of the scheme generated by grouping method increases fast. Recently, the authors in [9] proposed an interesting combinatorial structure called placement delivery array (PDA) to study the coded caching scheme with low subpacketization. They also showed that the MN scheme is equivalent to a special PDA which is called MN PDA. Based on the concept of PDA, there are many works focusing on achieving the low subpacketization, such as [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
@@
There are other characterizations of coded caching schemes, such as the linear block codes [21], the special $(6,3)$-free hypergraphs [22], the $(r,t)$ Ruzsa-Szeméredi graphs[23], the strong edge coloring of bipartite graphs[24], cross resolvable designs[25], the projective geometry [26], and combinatorial designs [27] which are listed in Table I. Here we do not include the schemes from in [23] and the first scheme in [28] in the Table I, since we focus on explicit constructions of schemes, while theirs only focus on the existing and the user number approximates infinite integer.
@@
Schemes & References Number of Users K Memory ratio Load Subpacketization Constrain
MN Scheme [2] $K$ $\frac{t}{K}$ $\frac{K-t}{t+1}$ (${K\choose t}$ $t\leq K$
WCLC scheme [12, 22, 19]
@@
YTCC scheme [24] $\binom{H}{a}$ $1-\frac{\binom{a}{r}\binom{H-a}{z-r}}{\binom{H}{z}}$
@@
WCCLS scheme [15] $q^{m}$ $1-\frac{\binom{m}{w}(q-1)^{w}}{q^{m}}$ $\frac{\binom{m}{w}(q-1)^{w}}{q^{m-w}}$ $q^{m}$ $m,w\in\mathbb{Z}^{+},m<w$
CKSM scheme 1 [26]
@@
XXGL scheme [28] $K$ $\frac{K-2}{K}$ $\frac{K-1}{K}$ $K$ $K\in\mathbb{N}^{+}$
AST scheme [20] $2^{r}k$ $1-\frac{r+1}{2^{r}}+\frac{r}{2^{r}k}$ $\frac{k(r+1)-r}{2^{r}}$ $2^{r}k=K$ $r,k\in\mathbb{N}^{+}$
MR scheme [29] $K$ $\frac{t}{K}$ $\left\lceil\frac{K(K-t)}{2+\left\lfloor\frac{t}{K-t+1}\right\rfloor+\left\lfloor\frac{t-1}{K-t+1}\right\rfloor}\right\rceil\cdot\frac{1}{K}$ $K$ $K,t\in\mathbb{Z}^{+},t<K$
By Table I, we can see that the schemes in [8, 2, 9, 10, 21, 22, 13, 19, 12] have flexible number of users, large memory regimes, and small loads, but large subpacketizations which increase exponentially (or sub-exponentially) with the number of users; the schemes in [24, 16, 14, 30, 20, 15, 26, 27, 28, 29] have low subpacketizations which increase polynomially or linearly with the number of users. However, the schemes in [26, 24, 15] have the special numbers of users (i.e., combinations, powers, or products of combinations and powers) and special memory ratio (i.e., the ratios of these combinations, powers, or products of combinations and powers); the schemes in [16, 27, 28] have the large memory ratio which approximates $1$; the schemes in [20, 26] have the memory ratios approximating 00 or $1$. The schemes in [29, 14] have the subpacketizations which increase linearly with the number of users for the flexible number of users and large memory regimes. It is worth noting that the schemes in [29, 14] use the same placement strategy, consecutive cyclic uncoded placement. The authors showed that under the consecutive cyclic uncoded placement and one-shot delivery, the maximum coded caching gain is $12\lfloor\frac{K}{K-KM/N+1}\rfloor+$ 1. Clearly when $M/N$ is small such that $M/N\leq 1/2$, the linear coded caching schemes have the maximum coded caching gain $3$.
@@ I-A Contribution
In this paper, we focus on constructing the coded caching scheme with linear subpacketization. When $K$ is odd, we introduce a new concept called Non-half-sum disjoint packing (NHSDP) which can realize a coded caching scheme with linear subpacketization. Compared to the existing characterizing method, the greatest advantage of our NHSDP is that it integrates the placement strategy and the transmission strategy into a single condition, i.e., the second condition of NHSDP in Definition 2. The main result can be summarized as follows.
@@
  • When $K$ is odd, we transform the coded caching construction to a new combinatorial structure named Non-half-sum disjoint packing (NHSDP). Given a $(v,g,b)$ NHSDP for any odd positive integers $v$, $g$, and $b$, we can obtain a $(K=v,M,N)$ coded caching scheme with the memory ratio $\frac{M}{N}=1-\frac{bg}{v}$, the subpacketization $F=K$, and the transmission load $R=b$. For the even number of users $K$, we can add one virtual user into the system such that the effective number of users is $K^{\prime}=K+1$, which could be solved by the $(K^{\prime}+1,g,b)$ NHSDP.
@@
  • By constructing an NHSDP, we can obtain a $(K=v,M,N)$ coded caching scheme with the memory ratio $M/N=1-\frac{2^{n}\prod_{i=1}^{n}m_{i}}{v}$, the coded caching gain $g=2^{n}$, and the transmission load $R=\prod_{i=1}^{n}m_{i}$ for any odd integer $v\geq 2\phi(m_{1},m_{2},\ldots,m_{n})+1$ where $n$ and $m_{1}$, $m_{2}$, $\ldots$, $m_{n}$ are any positive integers. In particular, when $m_{1}=\cdots=m_{n}=\lfloor\frac{v^{1/n}-1}{2}\rfloor$, we obtain a $(K=v,M,N)$ coded caching scheme with the memory ratio $\frac{M}{N}=1-(\frac{2\lfloor\frac{q-1}{2}\rfloor}{q})^{n}$, the subpacketzation $F=K$, the coded caching gain $g=2^{n}$, and the transmission load $R={\left\lfloor\frac{q-1}{2}\right\rfloor}^{n}$.
@@
  • Theoretical and numerical comparisons show that the proposed scheme achieves a lower load compared to the existing schemes with linear subpacketization (to the number of users); compared to some existing schemes with polynomial subpacketization, our proposed scheme can achieve a lower transmission load and also has lower load in some cases; compared to some existing schemes with exponential subpacketization, our scheme has loads close to those of these schemes in some cases.
  • The new concept of NHSDP has a close relationship between the classic combinatorial structures, such as cyclic difference packing (CDP) [31], non-three-term arithmetic progressions (NTAP) [32], and perfect hash family (PHF) [33], etc.
@@
    Specifically, a CDP can be used to construct a $(v,g,1)$ NHSDP. Given a $(v,g,1)$ NHSDP is equivalent to an NTAP set over $[v]$ and it can be used to construct a $(3:gv,v,3)$ PHF. In cryptography [33], combinatorics [34], and coding theory [35], we aim to construct a subset of $\mathbb{Z}_{n}$ satisfying NTAP, with the maximum cardinality, and a $(3:m,v,3)$ PHF with the maximum value of $m$. When $v=3^{n}$, we can obtain a class of NTAPs using our new NHSDP whose size is larger than the state-of-the-art achievable bound on the NTAP size in [36] when $n\leq 52$.
    The PHF derived from our proposed NHSDP has more columns compared to the first quadrics PHF[37]. Furthermore, the number of columns in our proposed PHF is close to that of the Hermitian PHF presented in [37].
@@ I-B Organizations and notations
The rest of this paper is organized as follows. In Section II, a coded caching caching system, concept of PDA and their relationship are introduced. In Section III, we introduce the concept of the non-half-sum disjoint packing (NHSDP) and show that it can be used to generate a PDA. In Section IV, we propose a new class of PDAs by constructing an NHSDP. In Section V, we provide the performance analysis of the proposed scheme and we expand the application of NHSDP in Section VI. Finally, we conclude this paper in Section VII.
Notation: In this paper, we will use the following notations. Let bold capital letter, bold lowercase letter and curlicue letter denote array, vector and set respectively; let $|\mathcal{A}|$ denote the cardinality of the set $\mathcal{A}$; define $[a]=\{1,2,\ldots,a\}$ and $[a:b]$ is the set $\{a,a+1,\dots,b-1,b\}$; $\lfloor a\rfloor$ denotes the largest integer not greater than $a$. $a\nmid b$ denotes $a$ does not divide $b$. Define that $\left[k\atop t\right]_{q}=\frac{(q^{k}-1)\dots(q^{k-t+1}-1)}{(q^{t}-1)\dots(q-1)}$, $\left\langle K\right\rangle_{t}=K\mod t$; $\mathbb{Z}_{v}$ is a positive integer ring.

### v3

# A New Construction Structure on Coded Caching with Linear Subpacketization: Non-Half-Sum Disjoint Packing

@@ Abstract
Coded caching is a promising technique to effectively reduce peak traffic by using local caches and the multicast gains generated by these local caches. Coded caching schemes have been widely investigated following the seminal work of Maddah-Ali and Niesen. Explicit coding constructions have been proposed for a variety of network topologies with information-theoretically optimal or near-optimal transmission load $R$. An important parameter in these constructions is the subpacketization $F$, i.e., the number of subpackets that each content file needs to be divided into. In particular, the original scheme of Maddah-Ali and Niesen as well as several other variants require $F$ to grow exponentially with the number of users $K$. In practice, files have finite size and too large $F$ yields impractically small subpackets. Therefore, it is important to design coded caching schemes with $F$ and $R$ as small as possible. At present, the few known schemes with subpacketization linear in $K$ achieve large load. In this paper, we consider the linear scaling regime $F=O(K)$ and design schemes with unprecedented transmission load $R$. Specifically, we first introduce a new combinatorial structure called non-half-sum disjoint packing (NHSDP) which can be used to generate a coded caching scheme with $K=O(F)$. A class of new schemes is then obtained by constructing NHSDP. Theoretical and numerical analyses demonstrate that (i) in comparison to existing schemes with linear subpacketization, the proposed scheme achieves a lower load; (ii) the proposed scheme also attains a lower load than some existing schemes with polynomial subpacketization in some cases; and (iii) the proposed scheme achieves load values comparable to those of existing schemes with exponential subpacketization in some cases. Furthermore, the newly introduced concept of NHSDP is closely related to classical combinatorial structures, including cyclic difference packings (CDP), non-three-term arithmetic progressions (NTAP), and perfect hash families (PHF). These relationships underscore the significance of NHSDP as an important combinatorial structure in the field of combinatorial design.
@@
Coded caching, placement delivery array, linear subpacketization, transmission load, non-half-sum disjoint packing
@@ I INTRODUCTION
The rise of internet connected devices in recent years has significantly increased network traffic, fueled by activities such as multimedia streaming, web browsing, and social networking. In addition, the high temporal variability of this traffic leads to congestion during peak periods and inefficient use of network resources during off-peak times. Caching is a promising technique to reduce peak traffic by taking advantage of memories distributed across the network to duplicate content during off-peak times. Conventional uncoded caching techniques focus on predicting the user demands to make appropriate prefetching decisions, thus realizing a “local caching gain”, which scales with the amount of local memory [1]. In their seminal paper [2], Maddah-Ali and Niesen (MN) demonstrated that, in addition to a local caching gain, coded caching can also attain a “global caching gain”. This global caching gain scales with the global amount of memory in the network, since each transmission of the MN scheme can serve multiple users simultaneously.
@@
Following [2], the basic coded caching problem consists of the following setup: a central server has $N$ equal-size files and it is connected to $K$ users via a shared (broadcast) link of normalized capacity, which is able to transmit one file per unit of time. Each user has a local cache memory of capacity equivalent to $M$ files. An $F$-division $(K,M,N)$ coded caching scheme consists of two phases: the placement phase during the off-peak hours and the delivery phase during the peak hours. In the placement phase, the server divides each file into $F$ equal-size packets and places some of these packets into the users’ caches without knowledge of the users’ future demands. If packets are directly placed in the user’s caches, the scheme is referred to as “uncoded placement”. If the server places more general functions of the packets in the users’ caches, the scheme is referred to as “coded placement”. The parameter $F$ is referred to as the subpacketization order of the scheme (in brief, subpacketization).
In the delivery phase, each user requests an arbitrary file. According to the users’ demands and the cache contents, the server broadcasts some coded packets such that each user can recover its desired file. The normalized amount of transmission for the worst-case over all possible demands is called the transmission load (or load) $R$. In particular, a scheme achieving load $R$ can serve all users’ demands by transmitting at most $R$ times the size of one file. Following [2], we define the caching gain as the reduction factor in the load with respect to sending sequentially all demanded files. For example, in the case of $K\leq N$ and distinct demands, a system without caching needs to send all $K$ files through the shared link. A system with traditional uncoded caching (e.g., using prefix caching, where a fraction $M/N$ of each file is cached) requires the transmission of $(1-M/N)K$ files. A coded caching system achieving load $R$ achieves a “coded caching gain” $K(1-M/N)/R\geq 1$ over the system without caches and the system with conventional uncoded caching.
@@
The first well-known coded caching scheme was proposed by Maddah-Ali and Niesen in [2], which utilized a combinatorial design in the placement phase and linear coding in the delivery phase. This scheme is referred to as the MN scheme. The authors in [3, 4, 5, 6] separately proved the minimality of the load of the MN scheme under the uncoded placement and for the case $K\leq N$. Using the placement strategy of the MN scheme, the authors in [5] proposed an exact optimal caching scheme, which strictly improves upon the current state of the art by exploiting commonalities among user demands under uncoded placement.
@@
However, the subpacketization $F={K\choose KM/N}$ of the MN scheme increases exponentially with the number of users $K$. This would become infeasible for practical implementation when $K$ is large. So it is meaningful to design a coded caching scheme with low subpacketization. The grouping method, which is widely regarded as the most effective in reducing subpacketization, was proposed in [7, 8]. Nevertheless, the load of the scheme generated by grouping method increases fast. Recently, the authors in [9] proposed an interesting combinatorial structure called placement delivery array (PDA) to study coded caching schemes with low subpacketization. They also showed that the MN scheme is equivalent to a special PDA which is called MN PDA. A large amount of articles have used the PDA approach to devise low subpacketizaton coded caching schemes (e.g., [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]).
@@
There are other characterizations of coded caching schemes, such as using linear block codes [21], the special $(6,3)$-free hypergraphs [22], the $(r,t)$ Ruzsa-Szeméredi graphs[23], the strong edge coloring of bipartite graphs[24], cross resolvable designs[25], the projective geometry [26], and combinatorial designs [27], which are listed in Table I. In this table, we do not include the schemes from [23] and the first scheme in [28], since we focus on explicit constructions, while theirs only focus on the existence of constructions and the user number approximates an infinite integer.
@@
Schemes Number of Users Memory ratio Load Subpacketization Constraint
MN Scheme [2] $K$ $\frac{t}{K}$ $\frac{K-t}{t+1}$ (${K\choose t}$
WCLC [19]
@@
YTCC [24] $\binom{H}{a}$ $1-\frac{\binom{a}{r}\binom{H-a}{z-r}}{\binom{H}{z}}$
@@
WCCLS [15] $q^{m}$ $1-\frac{\binom{m}{w}(q-1)^{w}}{q^{m}}$ $\frac{\binom{m}{w}(q-1)^{w}}{q^{m-w}}$ $q^{m}$ $m,w\in\mathbb{Z}^{+},m<w$
CKSM 1 [26]
@@
XXGL [28] $K$ $\frac{K-2}{K}$ $\frac{K-1}{K}$ $K$ $K\in\mathbb{N}^{+}$
AST [20] $2^{r}k$ $1-\frac{r+1}{2^{r}}+\frac{r}{2^{r}k}$ $\frac{k(r+1)-r}{2^{r}}$ $2^{r}k=K$ $r,k\in\mathbb{N}^{+}$
MR [29] $K$ $\frac{t}{K}$ $\left\lceil\frac{K(K-t)}{2+\left\lfloor\frac{t}{K-t+1}\right\rfloor+\left\lfloor\frac{t-1}{K-t+1}\right\rfloor}\right\rceil\cdot\frac{1}{K}$ $K$
From Table I, we notice that the schemes in [8, 2, 9, 10, 21, 22, 13, 19, 12] have a flexible number of users, large memory regimes, and a small load, but a large subpacketization which increases exponentially (or sub-exponentially) with the number of users; the schemes in [24, 16, 14, 30, 20, 15, 26, 27, 28, 29] have a low subpacketization which increases polynomially or linearly with the number of users. However, the schemes in [26, 24, 15] are restricted to the special number of users (i.e., combinations, powers, or products of combinations and powers) and special memory ratio (i.e., the ratios of these combinations, powers, or products of combinations and powers); the schemes in [16, 27, 28] have a large memory ratio close to $1$; the schemes in [20, 26] have memory ratios close to 00 or $1$. The schemes in [29, 14] have a large subpacketization increasing linearly with the number of users for the flexible number of users and large memory regimes. It is worth noting that the schemes in [29, 14] use the same placement strategy, namely, the so-called consecutive cyclic uncoded placement. The authors showed that under the consecutive cyclic uncoded placement and clique-covering delivery, the maximum coded caching gain is $12\lfloor\frac{K}{K-KM/N+1}\rfloor+$ 1. Clearly, when $M/N$ is small such that $M/N\leq 1/2$, linear coded caching schemes with one-short delivery can achieve a coded caching gain of at most $3$.
@@ I-A Contribution
In this paper, we focus on constructing coded caching schemes with linear subpacketization. When $K$ is odd, we introduce a new combinatorial concept called non-half-sum disjoint packing (NHSDP), which can realize a coded caching scheme with linear subpacketization. Compared to the existing characterizing methods, the greatest advantage of NHSDP is that it integrates the placement strategy and the transmission strategy into a single condition, i.e., the second condition of NHSDP in Definition 2. The main result can be summarized as follows.
@@
  • When $K$ is odd, we transform the coded caching construction to a new combinatorial structure named non-half-sum disjoint packing (NHSDP). Given a $(v,g,b)$ NHSDP for any odd positive integers $v$, $g$, and $b$, we can obtain a $(K=v,M,N)$ coded caching scheme with the memory ratio $\frac{M}{N}=1-\frac{bg}{v}$, subpacketization $F=K$, and transmission load $R=b$. If $K$ is even, we can add one virtual user to the system such that the effective number of users is $K^{\prime}=K+1$, and then construct a scheme based on the $(K’+1,g,b)$ NHSDP.
@@
  • By constructing an NHSDP, we can obtain a $(K=v,M,N)$ coded caching scheme with the memory ratio $M/N=1-\frac{2^{n}\prod_{i=1}^{n}m_{i}}{v}$, the coded caching gain $g=2^{n}$, and the transmission load $R=\prod_{i=1}^{n}m_{i}$ for any odd integer $v\geq 2\sum_{i=1}^{n-1}(m_{i}\prod_{j=i+1}^{n}(1+2m_{i}))+m_{n}+1$ where $n$ and $m_{1}$, $m_{2}$, $\ldots$, $m_{n}$ are any positive integers. In particular, when $m_{1}=\cdots=m_{n}=\lfloor\frac{v^{1/n}-1}{2}\rfloor$, we obtain a $(K=v,M,N)$ coded caching scheme with the memory ratio $\frac{M}{N}=1-\frac{2^{n}\lfloor\frac{v^{1/n}-1}{2}\rfloor^{n}}{v}$, subpacketization $F=K$, coded caching gain $g=2^{n}$, and transmission load $R={\left\lfloor\frac{v^{1/n}-1}{2}\right\rfloor}^{n}$.
@@
  • Theoretical and numerical comparisons show that our proposed scheme achieves a lower load compared to the existing schemes in [20, 14, 16, 28] with linear subpacketization; the proposed scheme achieves a lower load than some existing schemes in [26, 24] with polynomial subpacketization under some system parameters; compared to some existing schemes in [2, 19] with exponential subpacketization, our scheme incurs only a slight load increase for some system parameters.
@@
  • The new concept of NHSDP has a close relationship to other classic combinatorial structures, such as cyclic difference packing (CDP) [31], non-three-term arithmetic progressions (NTAP) [32], and perfect hash family (PHF) [33], etc. Specifically, a CDP can be used to construct a $(v,g,1)$ NHSDP. A $(v,g,1)$ NHSDP is equivalent to an NTAP set over $[v]$ and it can be used to construct a $(3:gv,v,3)$ PHF. In cryptography [33], combinatorics [34], and coding theory [35], we aim to construct a subset of $\mathbb{Z}_{n}$ satisfying NTAP, with the maximum cardinality, and a $(3:m,v,3)$ PHF with the maximum value of $m$. When $v=3^{n}$, we can obtain a class of NTAPs using our new NHSDP whose size is larger than the state-of-the-art achievable bound on the NTAP size in [36] when $n\leq 52$. The PHF derived from the proposed NHSDP has more columns compared to the first quadrics PHF[37]. Furthermore, the number of columns in the resulting PHF is close to that of the Hermitian PHF presented in [37].
@@ I-B Organizations and notations
The rest of this paper is organized as follows. In Section II, a coded caching system, the PDA, and their relationship are introduced. In Section III, we introduce the non-half-sum disjoint packing (NHSDP) and show that it can be used to generate a PDA. In Section IV, we propose a new class of PDAs by constructing an NHSDP. In Section V, we provide the performance analysis of the proposed scheme, and we expand the application of NHSDP in Section VI. Finally, we conclude this paper in Section VII.
Notation: In this paper, we will use the following notations. Let bold capital letter, bold lowercase letter, and curlicue letter denote array, vector, and set respectively; let $|\mathcal{A}|$ denote the cardinality of the set $\mathcal{A}$; define $[a]=\{1,2,\ldots,a\}$ and $[a:b]$ is the set $\{a,a+1,\dots,b-1,b\}$; $\lfloor a\rfloor$ denotes the largest integer not greater than $a$. $a\nmid b$ indicates that $a$ does not divide $b$. We define that $\left[k\atop t\right]_{q}=\frac{(q^{k}-1)\dots(q^{k-t+1}-1)}{(q^{t}-1)\dots(q-1)}$, $\left\langle K\right\rangle_{t}=K\mod t$; $\mathbb{Z}_{v}$ is the ring of integer residues modulo $v$.